OpenStudy (idealist10):

Find all curves y=y(x) such that the tangent to the curve at any point (xo, y(xo)) intersects the x axis at xI=(xo)^3.

OpenStudy (idealist10):

@myininaya

OpenStudy (anonymous):

The tangent line at any point \((x_0,y_0)\) to the curve \(y=y(x)\) has the general form \[y-y_0=y'(x_0)(x-x_0)\] You want to guarantee that the curves pass through the point \((x_i,0)=({x_0}^3,0)\), so plugging in \(x={x_0}^3\) and \(y=0\) into the tangent line equation gives the ODE, \[0-y_0=y'(x_0)({x_0}^3-x_0)\] which is linear in \(y_0=y(x_0)\), and also separable. Solve whichever way you like.

OpenStudy (idealist10):

Thanks!